OpenStudy (anonymous):

The hypotenuse of a right triangle is 50 millimeters long. One leg of the right triangle is 30 millimeters long. What is the length of the other leg? 20 millimeters 30 millimeters 40 millimeters 60 millimeters

OpenStudy (danjs):

pythagorean theorem

OpenStudy (anonymous):

what

OpenStudy (cheerleader):

you need to add all the sides up and they need to add up to 180 degrees

OpenStudy (anonymous):

\[a^2 + b^2 = hypotenuse^2\]

OpenStudy (anonymous):

ohhhhhhhh okay

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

No its not D

OpenStudy (danjs):

50^2 = 30^2 + x^2

OpenStudy (anonymous):

b?

OpenStudy (danjs):

2500 = 900 + x^2

OpenStudy (danjs):

x^2 = 1600

OpenStudy (danjs):

x = sqrt(1600) = 40

OpenStudy (anonymous):

thats 58.31 but that rounds to 60